Une the substitutive method to show that 7(n) = 27A) + n is in O(n). (You do not need to show that T(n) is in 2(n)
Answer #1

Answer:-

Given that CỦn) = 2 (7) +0 -0 Tha)= 270)-0 The equation and equation © are not part of substitution Now, guess Tin) - on; mea